OpenStudy (anonymous):

A solid cone of radius 6 cm and height 21 cm is melted and made into 2 spheres of different sizes.The radius of the first sphere is 4 cm , find the radius of the second sphere.

Miracrown (miracrown):

Hey. Lets do this, shall we? :)

OpenStudy (anonymous):

Yes !

Miracrown (miracrown):

This problem revolves around surface areas. First we need the surface area of a cone, do you know this formula?

OpenStudy (anonymous):

No :(

Miracrown (miracrown):

It's ok :) Here it is:

Miracrown (miracrown):

\[A = \pi r \space (r \space +\sqrt{h ^{2} \space + r^2} )\] \[\pi \space \times 6(6+\sqrt{21^2+6^2}\] = \[6\pi(6+\sqrt{477}) = 524.78\]

Miracrown (miracrown):

Tis our formula for the cone's surface area where r = radius and h = height if we use an approxmation for the area of the cone we should get about 524.7

OpenStudy (anonymous):

Ok!

Miracrown (miracrown):

Sorry, I'm in a hurry my cat just went berserk. Here I've done it for you in steps, please go through them and if you stumble upon anything then just shoot me a message and I'll be able to help you out once I get back. Take care! http://prntscr.com/857gfg